OpenStudy (crashonce):

cos(θ+15)−tan(θ−15)=(4cos2θ)÷(1+2sin2θ)

OpenStudy (crashonce):

@ikram002p pleas help me

OpenStudy (crashonce):

@ikram002p please

OpenStudy (crashonce):

@ganeshie8 @ikram002p @zepdrix sorry for extra tagging

OpenStudy (anonymous):

cos(theta - 15)) or there is division sign in between??

OpenStudy (anonymous):

Sorry, + is there..

OpenStudy (crashonce):

no its cos (blah) - tan (blah) = blah

OpenStudy (crashonce):

yea its ok c:

OpenStudy (shadowlegendx):

@ikram002p is having her BOYFRIEND!!! WHAT DOES THIS MEAN!??!?!?!?!?!??!

OpenStudy (crashonce):

bf = breakfast lolololol

OpenStudy (shadowlegendx):

SHE IS HAVING HER BOYFRIEND FOR BREAKFAST, OH MER GUSH :OOOOOOOOOOOOOOOOO

OpenStudy (crashonce):

ahhahahaahhacant argue with that but any pls

OpenStudy (crashonce):

pls help me on the original question

OpenStudy (anonymous):

I just want to know that if we can use the fraction values of cos(15) or sin(15) here???

OpenStudy (crashonce):

idk if u can

OpenStudy (aum):

The 15 degree angle is on the LHS but not on the RHS. Expand the LHS and get rid of the 15 degree angle by substituting (45-30) in its place since the trig values for 45 and 30 degrees are known.

OpenStudy (crashonce):

show me how that would wrk in the fraction bcos it doesnt seem t work

OpenStudy (aum):

\( \sin(15) = \sin(45-30) = \sin(45)\cos(30) - \cos(45)\sin(30) = \\ \sqrt{2}/2 * \sqrt{3}/2 - \sqrt{2}/2 * 1/2= \sqrt{2}/4 * (\sqrt{3} - 1) \)

OpenStudy (anonymous):

@aum It is just like using fraction values of sin(15) and cos(15)..

OpenStudy (crashonce):

@aum what about theat

OpenStudy (anonymous):

\[\sin(15^{\circ}) = \frac{\sqrt{3}-1}{2 \sqrt{2}}\] \[\cos(15^{\circ}) = \frac{\sqrt{3} + 1}{2 \sqrt{2}}\] And hence: \[\tan(15^{\circ}) = \frac{\sqrt{3} - 1}{\sqrt{3} + 1}\]

OpenStudy (anonymous):

My calculation is still going lengthy and it is not reaching anywhere significant.. :(

OpenStudy (anonymous):

One way I started by using tan(x) = sin(x)/cos(x).. But that will lead to: \[\frac{\sqrt{3} + 2\cos(2 \theta) - 8\sin(\theta - 15)}{8\cos(\theta - 15)}\]

OpenStudy (anonymous):

May be I have chosen wrong path to go with.. I try it again.. :)

OpenStudy (aum):

It may be worthwhile first to make sure the identity is true for say theta = 0.

OpenStudy (aum):

And I don't think the given identity is true for theta = 0.

OpenStudy (anonymous):

Yeah, it is not true for given value : \(\theta\)..

OpenStudy (anonymous):

\(\theta\) being 0..

OpenStudy (aum):

Instead of an identity, you can treat this as an equation instead and solve for theta using a graphing calculator.